The annual interest earned on two investments is $479. One investment is in a 2.3% tax-free annual simple interest account, and the other investment is in a 5.3% annual simple interest CD. The total amount invested is $13,000. How much is invested in each account?
Answer by richwmiller(17219)

Total amount of money invested. $13000
x+y=13000,
Total interest for the year for the two accounts is $479
0.023*x+0.053*y=479
x=13000-y
We substitute for x
0.023*(13000-y)+0.053*y=479
We multiply out
299.0-0.023y+0.053*y=479
We combine like terms.
0.03*y=180.0
Isolate y
y=6000.00 at 5.3%
x=13000-y
Calculate x
x=7000.00 at 2.3%
We check
0.023*7000.0+0.053*6000.0=479
161.0+318.0=479
479.0=479
Since this statement is TRUE and neither amount is negative then it is ok